# Fairness for the People, by the People: Minority Collective Action

### v1

@@ B.3 Success Bound With Label Error
This proof follows closely the proof of Theorem 5 by Hardt et al. [10]. We start under the assumption of an optimal Bayes classifier, setting $\epsilon=0$.
@@
B.4 Label Error With Better Representation
For the following we assume a similar setting as in Section˜B.2, visualised as a 2D distribution in Figure˜6. We are given the majority data, and tasked with labeling the minority data. Assume all labels are distributed equally $\mathbb{P}[Y=1]=\mathbb{P}[Y=-1]=\frac{1}{2}$. The minority features $X_{\text{min}}$ are distributed as $X_{\text{min}}\sim\mathcal{N}(y\mu_{\text{min}},\Sigma_{\text{min}})$ with $X_{\text{min}}\in\mathbb{R}^{d}$. The label $\hat{y}_{\text{1NN}}^{(n)}$ is predicted according to a 1NN classifier from $n$ majority samples $\mathcal{D}_{n}={(x_{i},y_{i})}_{i=0}^{n}$. Majority samples with $y=+1$ are distributed as $X_{+}\sim\mathcal{N}(\mu,\Sigma)$, and with $y=+1$ are distributed as $X_{-}\sim\mathcal{N}(-\mu,\Sigma)$.
@@ B.4 Label Error With Better Representation
Appendix C Technical Details
C.1 Datasets
@@
The Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) dataset contains the data of criminal defendants in Broward county sheriff’s office in Florida with the task of predicting the recidivism risk. The label in this dataset represents whether the person re-offended and the sensitive attribute is the race. We follow the same data cleaning and pre-processing as Alghamdi et al. [40].
@@
The Adult dataset [30] contains demographic features of US citizens and is tasked with predicting the income level of an individual. The label represents if the individual has income higher than \$50,000 and the sensitive attribute we use is the race. We follow the same data cleaning and pre-processing as Alghamdi et al. [40].
@@
The High School Longitudinal Study of 2009 (HSLS) [31] contains details of high-school students across the US and the task is to predict the academic success of the students. The label represents the exam score and the sensitive attribute is the race. We follow the same data cleaning and pre-processing as Alghamdi et al. [40].
@@
Ding et al. [32] offer different classification tasks derived by US census data. In our work we used the pre-defined task of predicting level of income denoted as ACSIncome, where the data is already pre-processed. The label represents if the individual has income higher than \$50,000 and the sensitive attribute is the race.
@@
The waterbirds dataset [33] contains images of landbirds and waterbirds super-imposed on either land or water backgrounds, with the task of classifying the image as of a landbird or a waterbird. The label represents the type of bird, and the sensitive attribute is whether the background is land or water. To obtain the features, we used the output of the penultimate layer of a pre-trained ResNet-18 network from PyTorch  [^1] https://pytorch.org/vision/main/models/generated/torchvision.models.resnet18.html. We report the results on those features as Waterbirds-Full. We also performed PCA (using scikit-learn) and kept the first 85 principal components which retain about 75% of the variance, and report the results of these components as Waterbirds-PCA.
@@
The CivilComments dataset [34] is a collection of text comments found on the internet, with the goal of training a classifier to fairly detect toxicity. For this paper, we modified the dataset to keep only the comments that include either christian or muslim (but not both), with a label 0 meaning toxic and 1 meaning safe. To obtain the features, we used the word embeddings given by Hugging Face’s bert-base-uncased model [^2] https://huggingface.co/google-bert/bert-base-uncased. We report the results on those features as CivilComments-Full. We also performed PCA (using scikit-learn) and kept the first 100 principal components which retain about 75% of the variance, and report the results of these components as CivilComments-PCA.
C.2 Training
@@
The probabilities for RB-prob were inferred by training scikit-learn’s histogram-based gradient boosting classification tree on the majority data with the default parameters, and using its predict_proba function. For LFR [8] we used the implementation in Holistic AI’s open source library  [^4] https://github.com/holistic-ai/holisticai with the default parameters. For FARE [24] we used the official implementation  [^5] https://github.com/eth-sri/fare with hyperparameters $\gamma=0.85$, $k=200$ and $n=100$. For all distance computation we used the Euclidean norm $\ell^{2}$-norm as $d\left(v,u\right)=\left\|v-u\right\|_{2}=\sqrt{\sum_{i}\left(v_{i}-u_{i}\right)^{2}}$.
Appendix D Additional Results
D.1 Comparison with prior work
We compare our method RB-prob with the existing methods KDP [7] and CND [6] in Figure˜10. Figure˜10 shows that our method, motivated by the counterfactual labeling, is more efficient in terms of required number of label flips, than the existing works.
@@ D.1 Comparison with prior work
Figure 10: Fairness per number of label flips of the Random baseline, our method RB-prob, and the existing methods KDP [7] and CND [6]. Our method is more efficient than prior work, requiring less flips to achieve the same level of fairness. Note that in this experiment CND could flip any label, while all other methods were restricted to the labels of 30% of the minority.
D.2 Expanded results
@@
Figure 14: The firm-side pre-processing method FARE [24] and the post-processing method calibrated equalized odds [36] attain 0 EqOd with large error, while RB-prob with $\alpha=0.3$ (Section˜3) has much smaller error and less unfairness than the base classifier, but unable to get 0 EqOd.
@@
Figure 15: The Pareto fronts for using a fair representation when computing the KNN for RB-dist dominate the Pareto fronts for KNN computed on untransformed features. The blue stars represent the KNN without transforming the data, and the yellow triangles represent the KNN when the data is transformed using FARE [24]. The lines are fitted by a polynomial of degree 2 to guide the eye.

### v2

@@ C.3 Success Bound With Label Error
This proof follows closely the proof of Theorem 5 by hardtAlgorithmicCollectiveAction2023. We start under the assumption of an optimal Bayes classifier, setting $\epsilon=0$.
@@
C.4 Label Error With Better Representation
For the following we assume a similar setting as in Section˜C.2, visualised as a 2D distribution in Figure˜6. We are given the majority data, and tasked with labeling the minority data. Assume all labels are distributed equally $\mathbb{P}[Y=1]=\mathbb{P}[Y=-1]=\frac{1}{2}$. The minority features $X_{\text{min}}$ are distributed as $X_{\text{min}}\sim\mathcal{N}(y\mu_{\text{min}},\Sigma_{\text{min}})$ with $X_{\text{min}}\in\mathbb{R}^{d}$. The label $\hat{y}_{\text{1NN}}^{(n)}$ is predicted according to a 1NN classifier from $n$ majority samples $\mathcal{D}_{n}={(x_{i},y_{i})}_{i=0}^{n}$. Majority samples with $y=+1$ are distributed as $X_{+}\sim\mathcal{N}(\mu,\Sigma)$, and with $y=+1$ are distributed as $X_{-}\sim\mathcal{N}(-\mu,\Sigma)$.
@@ C.4 Label Error With Better Representation
Appendix D Technical Details
D.1 Datasets
@@
The Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) dataset contains the data of criminal defendants in Broward county sheriff’s office in Florida with the task of predicting the recidivism risk. The label in this dataset represents whether the person re-offended and the sensitive attribute is the race. We follow the same data cleaning and pre-processing as alghamdiAdultCOMPASFair2022.
@@
The Adult dataset [beckerAdult1996] contains demographic features of US citizens and is tasked with predicting the income level of an individual. The label represents if the individual has income higher than \$50,000 and the sensitive attribute we use is the race. We follow the same data cleaning and pre-processing as alghamdiAdultCOMPASFair2022.
@@
The High School Longitudinal Study of 2009 (HSLS) [jeongFairnessImputationDecision2022] contains details of high-school students across the US and the task is to predict the academic success of the students. The label represents the exam score and the sensitive attribute is the race. We follow the same data cleaning and pre-processing as alghamdiAdultCOMPASFair2022.
@@
dingRetiringAdultNew2021 offer different classification tasks derived by US census data. In our work we used the pre-defined task of predicting level of income denoted as ACSIncome, where the data is already pre-processed. The label represents if the individual has income higher than \$50,000 and the sensitive attribute is the race.
@@
The waterbirds dataset [sagawaDistributionallyRobustNeural2020] contains images of landbirds and waterbirds super-imposed on either land or water backgrounds, with the task of classifying the image as of a landbird or a waterbird. The label represents the type of bird, and the sensitive attribute is whether the background is land or water. To obtain the features, we used the output of the penultimate layer of a pre-trained ResNet-18 network from PyTorch  [^1] https://pytorch.org/vision/main/models/generated/torchvision.models.resnet18.html. We report the results on those features as Waterbirds-Full. We also performed PCA (using scikit-learn) and kept the first 85 principal components which retain about 75% of the variance, and report the results of these components as Waterbirds-PCA.
@@
The CivilComments dataset [borkanNuancedMetricsMeasuring2019] is a collection of text comments found on the internet, with the goal of training a classifier to fairly detect toxicity. For this paper, we modified the dataset to keep only the comments that include either christian or muslim (but not both), with a label 0 meaning toxic and 1 meaning safe. To obtain the features, we used the word embeddings given by Hugging Face’s bert-base-uncased model [^2] https://huggingface.co/google-bert/bert-base-uncased. We report the results on those features as CivilComments-Full. We also performed PCA (using scikit-learn) and kept the first 100 principal components which retain about 75% of the variance, and report the results of these components as CivilComments-PCA.
D.2 Training
@@
The probabilities for RB-prob were inferred by training scikit-learn’s histogram-based gradient boosting classification tree on the majority data with the default parameters, and using its predict_proba function. For LFR [zemelLearningFairRepresentations2013] we used the implementation in Holistic AI’s open source library  [^4] https://github.com/holistic-ai/holisticai with the default parameters. For FARE [jovanovicFAREProvablyFair2023] we used the official implementation  [^5] https://github.com/eth-sri/fare with hyperparameters $\gamma=0.85$, $k=200$ and $n=100$. For all distance computation we used the Euclidean norm $\ell^{2}$-norm as $d\left(v,u\right)=\left\|v-u\right\|_{2}=\sqrt{\sum_{i}\left(v_{i}-u_{i}\right)^{2}}$.
Appendix E Additional Results
E.1 Comparison with prior work
We compare our method RB-prob with the existing methods KDP [luongKNNImplementationSituation2011] and CND [kamiranClassifyingDiscriminating2009] in Figure˜10. Figure˜10 shows that our method, motivated by the counterfactual labeling, is more efficient in terms of required number of label flips, than the existing works.
@@ E.1 Comparison with prior work
Figure 10: Fairness per number of label flips of the Random baseline, our method RB-prob, and the existing methods KDP [luongKNNImplementationSituation2011] and CND [kamiranClassifyingDiscriminating2009]. Our method is more efficient than prior work, requiring less flips to achieve the same level of fairness. Note that in this experiment CND could flip any label, while all other methods were restricted to the labels of 30% of the minority.
E.2 Expanded results
@@
Figure 14: The firm-side pre-processing method FARE [jovanovicFAREProvablyFair2023] and the post-processing method calibrated equalized odds [pleissFairnessCalibration2017] attain 0 EqOd with large error, while RB-prob with $\alpha=0.3$ (Section˜3) has much smaller error and less unfairness than the base classifier, but unable to get 0 EqOd.
@@
Figure 15: The Pareto fronts for using a fair representation when computing the KNN for RB-dist dominate the Pareto fronts for KNN computed on untransformed features. The blue stars represent the KNN without transforming the data, and the yellow triangles represent the KNN when the data is transformed using FARE [jovanovicFAREProvablyFair2023]. The lines are fitted by a polynomial of degree 2 to guide the eye.